OpenStudy (anonymous):

What is the sum of a 6–term geometric series if the first term is 21 and the last term is 1,240,029?

OpenStudy (anonymous):

anyone? Please help me!

OpenStudy (amistre64):

do you recall any formulas that might help out?

OpenStudy (anonymous):

a(1)-a(a)r^n/1-r ?

OpenStudy (amistre64):

the formula for a sum of a geo progression requires us to find the first term, the number of terms, and the common ratio thats the one

OpenStudy (anonymous):

how do you find the common ratio though?

OpenStudy (amistre64):

and the formula for a general term of the a geo seq is?

OpenStudy (amistre64):

an = a1 * r^n (an/a1)^(1/n) = r

OpenStudy (amistre64):

err, make that n-1

OpenStudy (amistre64):

an = a1 * r^(n-1) (an/a1)^(1/(n-1)) = r better :)

OpenStudy (anonymous):

sorta. im still confused though. So wait I plug the numbers in and get 1,240,029/21^1/6-1

OpenStudy (amistre64):

a1=21, a6=that big number, and n=6\[r=\left(\frac{big}{21}\right)^{1/6}\]

OpenStudy (amistre64):

err, 1/5 yes

OpenStudy (amistre64):

im only slightly confused lol

OpenStudy (anonymous):

haha Im lost. this stuff confuses me more than anything

OpenStudy (amistre64):

so: 21(1-( 1240029/21)^6/5)/(1-(1240029/21)) is what i would type into the wolf for simplicity

OpenStudy (anonymous):

whats the wolf?

OpenStudy (amistre64):

thisis the wolf, and i syntaxed it wrong, this ones better http://www.wolframalpha.com/input/?i=21%281-%28+1240029%2F21%29%5E%286%2F5%29%29%2F%281-%281240029%2F21%29%29

OpenStudy (anonymous):

ughh thankyou but thats not an answer choice :( these are the choices 1,395,030 1,461,460 1,527,890 1,594,320

OpenStudy (amistre64):

lets chk my mathing then ....

OpenStudy (anonymous):

okay thankyou! I wish I understood this

OpenStudy (amistre64):

need the "r" value; everything else is given \[a_n=a_1*r^{n-1}\] \[\frac{a_n}{a_1}=r^{n-1}\] \[\left(\frac{a_n}{a_1}\right)^\frac{1}{n-1}=r\] \[\left(\frac{a_6}{a_1}\right)^\frac{1}{5}=r=\left(\frac{1240029}{21}\right)^\frac{1}{5}\]

OpenStudy (amistre64):

ohhh, i left off the 1/5 in the wolf lol

OpenStudy (anonymous):

oh my gosh! Thankyou so much! I get it now! See I didnt realize you had to use the same formula and solve for "r"! I thought it was a different formula that I didnt have! thanky so much!

OpenStudy (amistre64):

good luck :)